OpenStudy (anonymous):

find the exact value of x=cos(pi/12)

OpenStudy (campbell_st):

so you need to difference of 2 angles \[\frac{\pi}{4} - \frac{\pi}{6} = \frac{\pi}{12}\] then using the difference of 2 angles in cos you can use \[\cos(\frac{\pi}{4} - \frac{\pi}{6}) = \cos(\frac{\pi}{4})\times \cos(\frac{\pi}{6}) + \sin(\frac{\pi}{4}) \times \sin(\frac{\pi}{6})\] now you can substitute exact values for the ratios. hope it helps

OpenStudy (anonymous):

So the answer would be \[\sqrt{2}/2\]?

OpenStudy (anonymous):

Thank you!

OpenStudy (campbell_st):

no....  and